# Unmasking Deepfakes: Leveraging Augmentations and Features Variability for Deepfake Speech Detection

### v1

@@ 4 Analysis of Model Performance Across Deepfake Attacks
Examining detection difficulty across codecs revealed that C4 and C7 presented particular challenges, yielding consistently high EERs across models. This suggests that these codecs may mask or distort essential audio features that models rely on. Conversely, C2, C5, C6, and None (uncompressed) audio exhibited lower EERs, indicating that models were generally able to process and identify deepfake more effectively within these formats. This highlights the critical role of codec selection in the design and training of deepfake speech detection models. The results presented in Table 8 reveal that models utilizing $MS_{B}$ augmentations consistently outperform those trained with $MF_{G}$, indicating that $MS_{B}$ contributes to enhanced robustness across different codecs. While XLS-R 53 and XLS-R 128 generally achieve superior performance compared to ASVSpoof5+, the latter demonstrates notable effectiveness on C4 and C7, underscoring its resilience in challenging scenarios.
The fusion of all single models results in a more robust overall performance, significantly improving upon the results of single models for most codecs. However, for the most challenging codecs, C4 and C7, the fusion achieves good results but does not outperform one or more of the single models. The fusion of ASVSpoof5+ models consistently yields better performance than any individual ASVSpoof5+ model, demonstrating that combining models trained with $MS_{B}$ augmentation and those trained with $MF_{G}$ augmentation provides different perspectives on the data, resulting in a strong overall fusion performance.
@@ 5 Discussion
5.1 MaskedSpec vs MaskedFeature
Our results insights into the role of feature and spectral augmentations in deepfake speech detection. Feature-based augmentations (MaskedFeature) and raw-audio-based augmentations (MaskedSpec) each play distinct roles in model robustness. MaskedFeature focuses on perturbing the extracted feature space, forcing the model to adapt to variations in learned representations. MaskedSpec affects the raw audio, directly altering the signal’s spectral and temporal characteristics, thereby exposing the model to a broader range of distortions. The comparison between MaskedSpec (Table 2) and MaskedFeature (Table 4) highlights distinct impacts when applying the same masking strategies to different input representations. In MaskedSpec, where masks are applied directly to the audio input, Bands mask show lowest Eval EER values of 4.53%. Conversely, MaskedFeature masking on the feature-level representations shows a notable shift: the Gauss mask achieves the lowest Eval EER at 5.22%, while the Bands mask, effective in MaskedSpec, now shows increased Eval EER (8.78%). This suggests that masking at the feature representation level (MaskedFeature) introduces complexities different from those encountered in direct audio-level augmentation (MaskedSpec).
5.2 Pretrained Feature Extractor
The results presented in Table 5 emphasize the substantial influence of the pretrained feature extractor on the model’s performance. This is further evident in the fusion results, where models with different pretraining weights outperformed individual models, highlighting the complementary nature of diverse pretrained representations. Notably, while XLS-R 128 was trained on a larger and more linguistically diverse dataset, XLS-R 53 surprisingly achieved a lower Equal Error Rate (EER) during evaluation. This result suggests a more favorable compatibility of XLS-R 53 with the ASVSpoof5 evaluation dataset. In contrast, the Large-960h model, pretrained exclusively on the English-only Librispeech dataset, underperformed relative to the multilingual models. Given that the evaluation dataset is also English-only, these results suggest that multilingual pretraining improves the adaptability of the model and enhances the feature extractor’s ability to generalize.
Pretraining on the original ASVSpoof5 dataset, despite its task-specific focus on English and spoofed audio, surprisingly yielded a high EER. This outcome suggests a potential mismatch between the pretrained feature extractor and the requirements of the evaluation dataset. A plausible explanation is that the small size and limited diversity of the ASVSpoof5 dataset restrict the generalization capabilities of the pretrained model, which are critical for effective self-supervised learning.
When compression augmentations are introduced in the ASVSpoof5+ configuration, the evaluation EER improves significantly. This improvement suggests that the suboptimal performance observed with the ASVSpoof5 pretraining was likely due to the limited size and lack of variability in the dataset—factors critical for effective self-supervised learning. The introduction of compression augmentations enhances the model’s ability to generalize, thereby mitigating the challenges posed by the dataset’s constraints.
5.3 Mixed Augmentations
Mixed augmentations introduce variability that originally design to support generalization across diverse conditions, though excessive disruption can obscure key features and complicate learning. Table 3 illustrates that LPF and feature normalization generally improve Dev EER across mask types, but their impact on Eval EER is less consistent. For instance, Gauss masking with feature normalization achieves a low Dev EER (0.94%) but shows a modest Eval EER (5.82%), suggesting careful augmentations might support effective training metrics without assured generalization.
While feature normalization seems more beneficial with Gauss mask, LPF may better support simpler masks like Bands and Squares, likely by minimizing the disruption of essential features. These observations suggest that well-selected augmentations could enhance robustness, though the alignment between Dev and Eval metrics remains variable.
6 Conclusions and Future Work
This work introduces a deepfake speech detection model with a hybrid architecture, combining feature extractor architecture and a classifier head, trained end-to-end on the ASVSpoof5 dataset set. Despite limitations in dataset size, our model incorporates data augmentation and training strategies to achieve robust performance. Augmentation techniques applied at audio and feature levels have been shown to impact model performance in distinct ways, enhancing the variability within the training process. The MaskedSpec and MaskedFeature augmentation strategies, specifically utilizing the Bands and Gauss masks, not only enhance the model’s robustness but also significantly strengthen its performance in fusion configurations. These augmentations enable the model to generalize better across diverse conditions, contributing to improved overall effectiveness in handling complex synthetic audio scenarios. The results indicate that different mask shapes are well-suited to address various types of speech distortions.
Our findings indicate that multilingual pretraining, as utilized in the XLS-R 128 and XLS-R 53 models, enable our model to achieve lower evaluation EERs compared to monolingual models. Direct pretraining on ASVSpoof5 with the inclusion of compression augmentations (ASVSpoof5+) demonstrates strong performance, indicating that incorporating compression augmentations during pretraining is particularly effective when working with small, single-language datasets.
Notably, our model’s end-to-end training approach and augmentations establish a SoTA results for ASVSpoof5 benchmark in closed-set conditions and achieve even lower EERs when evaluated without closed conditions. Extension of this work can further explore pretraining augmentations and methods, as our findings suggest that different datasets capture diverse audio features. Extending augmentation strategies, such as incorporating masking augmentations directly into the pretraining phase, may further enhance the model’s robustness to distortions commonly introduced by spoofing techniques. Additionally, evaluating alternative classifier heads, with a focus on reducing model size while maintaining or improving performance, could lead to a more efficient model suitable for real-time applications.

### v2

@@ 4.2 Codecs
Examining detection difficulty across codecs revealed that C04 and C07 presented particular challenges, yielding consistently high EERs across models. This suggests that these codecs may mask or distort essential audio features that models rely on. Conversely, C01, C02, C05, C06, C11, and None (uncompressed) audio exhibited lower EERs, indicating that models were generally able to process and identify fake audio more effectively within these formats.
The results presented in Table 11 reveal that models pretrained on large and diverse corpora (XLS-R 128 and WavLM) consistently achieve lower EERs. Augmentation with MaskedSpec ($MS_{B}$) and compression ($C$) proves particularly effective across a wide range of codecs. For example, the XLS-R 128 model with $MS_{B}$ + C achieves the best performance on codecs such as None, C01, C06, C09, and C11, and maintains competitive results elsewhere.
Despite its limited data scale, the ASVSpoof5+ demonstrates notable improvements. While it underperforms compared to large-scale pretrained models on average, it achieves the lowest EERs in certain challenging codecs where XLS-R 128 struggles, such as C04 and C07. This indicates that task-specific distortions seen during ASVSpoof5+ pretraining may offer benefits for generalization under similar conditions.
4.3 Analysis-Driven Fusion Strategy
The analysis reveals that several models exhibit complementary strengths across different spoofing attacks and codec conditions. Notably, the models XLS-R 128 with $MS_{B}$ + $R$ + $C$, XLS-R 53 with $MS_{B}$ + $C$, and ASVSpoof5+ with both $MS_{B}$ + $C$ and $MF_{G}$ + $C$ demonstrate strong yet distinct performance profiles.
Although none of these models individually achieved the best performance on either the development or evaluation sets, their fusion leads to the lowest observed EER of 2.58% (Table 12). This highlights a critical insight: the effectiveness of model fusion lies not in combining top-performing models alone, but in leveraging the complementary error patterns and inductive biases across architectures and augmentations. Fusion strategies motivated by analytical diversity rather than raw accuracy can yield more robust and generalizable systems.
Fused Models ASVSpoof5 Eval EER (%)
XLS-R 128, $MS_{B}$ + $R$ + $C$
XLS-R 53, $MS_{B}$ + $C$
ASVSpoof5+, $MS_{B}$ + $C$
ASVSpoof5+, $MF_{G}$ + $C$ 2.58
Table 12: Fusion of four strategically selected models on the ASVspoof5 evaluation set, chosen for their complementary performance under codec and spoofing attack conditions. Although none are the top individual performers, their diverse strengths yield the lowest overall evaluation EER, demonstrating the benefit of cross-extractor, cross-augmentation fusion.
@@ 5 Discussion
5.1 Spectral and Feature-Level Masking
We evaluate four masking strategies at both the spectral and feature levels. The comparison of mask shapes within MaskedSpec and MaskedFeature (Tables 2 and 4) identifies the most effective configurations: Bands for MaskedSpec and Gauss for MaskedFeature.
From a theoretical perspective, these masking strategies can be viewed as structured forms of dropout, tailored to the audio domain. Spectral Bands masking acts as dropout in the frequency domain, occluding contiguous frequency regions and compelling the model to infer missing spectral content from global context rather than relying on narrowband artifacts. Feature-level Gauss masking operates within the latent representation space, where attenuating localized activations while preserving global structure promotes the development of distributed, noise tolerant embeddings. This effect mirrors observations in masked image modeling, where structured masking has been shown to encourage context-aware and generalizable representations.
These strategies provide complementary benefits to RawBoost and compression augmentations (Table 5), and their joint use yields consistent improvements, as reflected in the ASVSpoof5+ fusion results (Table 8). Importantly, the effectiveness of these masking methods remains stable across different pretraining conditions, highlighting their robustness and general applicability.
5.2 Augmentation in Self-Supervised Pretraining
To enhance representation learning under limited-domain conditions, we applied compression augmentation during the self-supervised pretraining phase of ASVSpoof5+. Our goal was to encourage the development of invariant features capable of generalizing across degraded acoustic conditions. The effect of this augmentation was substantial and led to over 60% relative improvment.
Interestingly, models trained with RawBoost during the end-to-end phase, but without exposure to it in pretraining, consistently underperformed. The RawBoost-augmented ASVSpoof5+ configurations yielded poor results and were ultimately excluded from our fusion ensemble. This suggests that certain augmentations may need to be introduced earlier in the training pipeline to fully integrate their distortive priors into the learned representation.
Notably, despite its smaller scale and limited linguistic diversity, the ASVSpoof5+ initialization was able to match or exceed the performance of much larger, general-purpose initializations such as XLS-R 128 and WavLM. This highlights the effectiveness of domain-specific augmentation strategies in improving self-supervised learning under constrained data regimes.
5.3 Fusion of Diverse Models and Augmentation Strategies
Another key outcome of our study is the benefit of combining models trained with different pretraining initializations and augmentation strategies. Pretrained feature extractors such as XLS-R 128 and ASVSpoof5+ encode speech representations shaped by the scale, language diversity, and domain specificity of their respective training corpora. Moreover, our augmentation techniques, includes augmentation applied at both the raw audio and feature levels, introduce varied perturbations that promote resilience to distributional shifts and unseen spoofing attacks.
Fusion proved to be an effective strategy for enhancing system robustness, resulting in improved performance under previously unseen distortions and manipulations. As demonstrated in our fusion experiments, combining models trained with both raw audio and feature-level augmentations, as well as using diverse pretraining configurations, reduced the evaluation EER to 2.81%, significantly outperforming any individual model. These results highlight the importance of leveraging complementary model characteristics, indicating that diversity in training settings contributes meaningfully to the development of resilient and generalizable systems.
6 Conclusion and Future Work
This study presents a comprehensive framework for deepfake speech detection that integrates a multi-stage augmentation strategy with hybrid training. We propose novel masking augmentations applied at both the raw audio spectrogram and learned representations, and conduct an empirical analysis of their effects across different masks shapes. Our findings indicate that Bands masking on spectrograms and Gauss masking on feature level are particularly effective, leading to consistent reductions in error rates across multiple model backbones.
Additionally, we introduce a targeted compression augmentation strategy applied during the self-supervised pretraining phase, addressing the vulnerability of self-supervision in low-resource scenarios. Introducing this augmentation resulted in a reduction of over 60% in EER and contributed to improved convergence during end-to-end training.
Our hybrid training framework combines self-supervised pretraining of the feature extractor with a ResNet-34 classification head, which are subsequently trained jointly end-to-end. On the ASVspoof5 (Track 1) benchmark, our system achieved state-of-the-art performance under the closed condition. Furthermore, when evaluated on ASVspoof2019 and ASVspoof2021 DF, the framework reached 0.18% and 2.92% EER respectively.
While our approach achieves leading results on multiple benchmarks, we believe future work should focus on creating cross-domain unified feature representations to further strengthen performance in real-world scenarios. Improving generalization across diverse audio sources remains a key challenge, and advancing feature extractors that learn consistent, domain-invariant representations will be critical for deploying reliable deepfake detection systems in practice.
Appendix A Full results with minDCF Results
While the official metric for the ASVSpoof5 Challenge is minDCF, our main paper reports EER due to its widespread use for deepfake speech detection. For completeness, we report minDCF in this appendix for multiple models, using the ASVSpoof5 cost configuration: $C_{\mathrm{miss}}=1$, $C_{\mathrm{fa}}=10$, and $\pi_{\mathrm{spoof}}=0.05$.
Pretraining Augmentations EER Dev (%) EER Eval (%) minDCF Eval
XLS-R 53 $MS_{B}$ + C 1.36 4.53 0.127
$MS_{B}$ + R 2.29 4.82 0.130
$MS_{B}$ + C + R 2.86 4.99 0.139
$MF_{G}$ + C 1.38 5.22 0.148
$MF_{G}$ + R 3.26 5.21 0.146
$MF_{G}$ + C + R 1.88 4.94 0.138
XLS-R 128 $MS_{B}$ + C 0.34 4.75 0.137
$MS_{B}$ + R 1.82 4.17 0.119
$MS_{B}$ + C + R 2.39 4.22 0.121
$MF_{G}$ + C 0.29 5.01 0.144
$MF_{G}$ + R 1.97 4.86 0.139
$MF_{G}$ + C + R 1.65 4.31 0.121
WAVLM $MS_{B}$ + C 0.86 4.77 0.132
$MS_{B}$ + R 3.64 5.54 0.156
$MS_{B}$ + C + R 1.69 5.40 0.154
$MF_{G}$ + C 2.57 6.96 0.188
$MF_{G}$ + R 3.62 5.51 0.153
$MF_{G}$ + C + R 2.44 5.38 0.154
ASVSpoof5+ $MS_{B}$ + C 5.12 5.44 0.117
$MS_{B}$ + R 17.20 12.23 0.334
$MS_{B}$ + C + R 17.87 14.89 0.384
$MF_{G}$ + C 5.62 5.27 0.144
$MF_{G}$ + R 17.23 10.25 0.249
$MF_{G}$ + C + R 20.33 12.90 0.299
Table 13: EER and minDCF results across different pretraining backbones and augmentations.
In table 13 we can see that EER and minDCF generally follow similar trends across models, indicating consistent relative performance. However, a notable exception emerges with ASVSpoof5+ trained with $MS_{B}$ + $C$, which achieves the lowest minDCF (0.117) but ranks outside the top ten in terms of EER (5.44%). This discrepancy highlights how minDCF, influenced by decision thresholds and cost priors, can favor models with better-calibrated outputs over those with the lowest raw error rates.